OpenStudy (anonymous):

Use order of operations to find a solution to the expression below. The solution to the expression is .

OpenStudy (anonymous):

Order of Operations. Acronym is PEMDAS

OpenStudy (anonymous):

Solve what's in the Parentheses first

OpenStudy (anonymous):

5 + (6 x [18/3 x 2] + 4) - 7 18/3 = 6 6 * 2 = 12 5 + (6 x 12 + 4} - 7 6 * 12 = 72 72 + 4 = 76 5 + 76 - 7 76 + 5 = 81 81 - 7 = 74 I believe that's correct

OpenStudy (anonymous):

click the link lol

OpenStudy (anonymous):

WAIT WAIT WAIT omg sorry I did division before multiplication and may have messed it all up

OpenStudy (anonymous):

Do you have answer choices?

OpenStudy (anonymous):

5 + (6 x [18/3 x 2] + 4) - 7 3*2 = 6 18/6 = 3 6 * 3 = 18 18 + 4 = 22 22 + 5 = 27 27 - 7 = 20 dang it sorry this should be right. i messed up

OpenStudy (anonymous):

The answer is 20

OpenStudy (anonymous):

i feel bad, i mislead that kid

OpenStudy (anonymous):

LOL at least you resolved it

OpenStudy (anonymous):

lol true, though he's not responding. he may have submitted it and is now rolling on the floor in tears because i failed him

OpenStudy (anonymous):

Hahahahahaha I can't

OpenStudy (anonymous):

their future down the drain all because i didn't freakin pay attention to PEMDAS